OpenStudy (sleepyjess):

Find the vertices and foci of the hyperbola with equation \(\dfrac{x^2}{16} - \dfrac{y^2}{48} = 1\) Vertices: (0, +- 4); Foci: (0, +- 8) Vertices: (0, +- 8); Foci: (0, +- 4) Vertices: (+- 4, 0); Foci: (+- 8, 0) Vertices: (+- 8, 0); Foci: (+- 4, 0)

OpenStudy (sleepyjess):

@ganeshie8

ganeshie8 (ganeshie8):

straightforward question how are you stuck on this ?

OpenStudy (sleepyjess):

I know the foci is +-4, but not sure how to find the vertices. \(c^2 = \sqrt{a^2 + b^2}\\c^2 = \sqrt{16+48}\\c^2 = \sqrt{64}\\c = 4\)

ganeshie8 (ganeshie8):

\[c^2 = a^2 + b^2\] you should get \(c = \pm 8\)

OpenStudy (sleepyjess):

Oh, I read it wrong, \(c = \sqrt{a^2 + b^2}\\c = \sqrt{16+48}\\c = \sqrt{64}\\c = 8\)

ganeshie8 (ganeshie8):